I cannot figure out this question:
The pressure at a certain depth in the ocean can be approximated by the equation P=15 + 1/2 D, where P is the pressure in pounds per square inch and D is the depth in feet. Use the equation to find the depth when the pressure is 40 lb/in 2.
isbn: 0618372628

Answer by stanbon(48558)

P=15 + 1/2 D
40=15 + (1/2)D
35=(1/2)D
D=70 ft.
Cheers,
Stan H.